# CrackSegFlow: Controllable Flow-Matching Synthesis for Generalizable Crack Segmentation with the CSF-50K Benchmark

### v1

@@ Abstract
Automated crack segmentation is essential for scalable condition assessment of pavements and civil infrastructure, yet practical deployment is limited by scarce pixel-level labels and severe domain shift across sensors, illumination, textures, and annotation conventions. This paper presents CrackSegFlow, a controllable flow-matching synthesis framework that generates photorealistic crack images conditioned on binary masks while preserving strict mask–image alignment. The generator combines topology-preserving mask injection with boundary-gated modulation to maintain thin-structure continuity and suppress texture-driven false positives. A second class-conditional flow-matching model synthesizes crack masks with explicit control over crack coverage, enabling balanced, topology-diverse paired data without additional manual annotation. We further inject crack masks into crack-free backgrounds to diversify illumination and surface artifacts and reduce false positives caused by shadows, joints, and pavement markings. Experiments on five benchmarks spanning four asphalt datasets and the crack class of a concrete-domain dataset demonstrate consistent improvements under an established hybrid CNN–Transformer segmentation backbone and a fixed training protocol. With real plus synthesized pairs, in-domain performance improves on average by 5.37 mIoU and 5.13 F1, and target-guided cross-domain synthesis yields average gains of 13.12 mIoU and 14.82 F1 using only limited target mask statistics. Compared with diffusion-based semantic synthesis, CrackSegFlow provides substantially faster deterministic sampling and improves fidelity and mask–image alignment for thin-structure crack geometry. Finally, we release CSF-50K, a public dataset of 50,000 paired crack images and pixel-accurate masks for large-scale benchmarking of generalizable crack segmentation.
@@ 1 Introduction
    We introduce CrackSegFlow, a controllable flow-matching (FM) synthesis framework that generates paired crack images and pixel-accurate masks with strict mask–image alignment. The image renderer integrates topology-preserving mask injection and boundary-gated modulation to preserve sub-pixel continuity of thin cracks and suppress texture-driven false positives under deterministic ODE sampling.
@@ 2 Related Works
Flow matching (FM) Lipman et al. [2022] and rectified flow Liu et al. [2022] generate by learning a time-dependent vector field that transports a simple base distribution to data along a prescribed path, enabling deterministic ODE sampling. Training regresses closed-form pairwise velocities along the path, which yields stable objectives and few-step sampling. In dense prediction, two directions have emerged. First, unified rectified-flow models bind image and mask spaces so a single network can both synthesize images and predict segmentation maps Wang et al. [2024a], Caetano and others [2025]. Second, segmentor-centric FM variants predict masks from images using flow-based objectives Bogensperger et al. [2025]. These works demonstrate FM’s viability but differ from an external, controllable augmentation engine; prior flow-based designs typically do not parameterize crack topology (e.g., explicit sparsity classes) nor guarantee geometry-consistent image rendering from masks at inference without additional labels.
@@ 3.2 Preliminaries
3.3 Semantic Flow-Matching Model
@@ 3.5 Rectified Flow for Background-Guided Crack Injection
Given a new background $b$ and mask $M$, we solve the deterministic ODE $\frac{dx}{d\tau}=v_{\theta}(x,\tau,M),\quad x(0)=b,\ \tau\in[0,1],$ with a fixed-step Euler integrator (e.g., $K$ uniform steps). The terminal state $x(1)$ preserves the background context while injecting the requested crack topology specified by $M$. This complements our standard flow-matching formulation by using the rectified schedule to “straighten” transport from background to target, yielding noise-free, deterministic sampling.
@@ 4.2.4 Synthesis protocol
As shown in  Fig.˜9, CrackSegFlow generates images that are precisely aligned with the conditioning masks. To maximize mask diversity, we first sample semantic masks using the mask generator and then render images conditioned on these masks. As depicted in  Fig.˜10, the resulting images are novel in crack geometry and texture relative to the originals. Table˜3 summarizes generative quality for both modalities. Across datasets, flow-matching achieves low FID/KID, with masks consistently scoring best (reflecting sharp topology), and images remaining competitive. These scores corroborate the visual fidelity in Figs. 9 to 10 support the downstream gains, where higher-quality pairs correlate with improved mIoU/F1 after augmentation across all datasets.
@@
Table 3: Quality of flow-matching synthesis measured by FID/KID (lower is better) for both mask and image modalities.
@@ 4.5 Comparison to diffusion-based semantic synthesis
Latent-diffusion pipelines, including LDM Rombach et al. [2022] and its conditioned variants such as ControlNet Zhang et al. [2023] and T2I-Adapter Mou et al. [2024], inject mask information through concatenation or auxiliary conditioning branches in the latent U-Net. While effective for general semantic layouts, this design can be sub-optimal for geometry-sensitive structures such as hairline cracks. In particular, repeated normalization and feature mixing can attenuate the conditioning signal at fine scales, leading to blurred filaments, weakened connectivity at junctions, and occasional mask–image drift Lei et al. [2025]. Fig.˜16 illustrates this limitation; latent-diffusion outputs often blur thin cracks or misalign them with the conditioning mask, whereas our Flow-Matching formulation replaces stochastic noise injection with a continuous, condition-aware transport that preserves topology and alignment. CrackSegFlow avoids this failure mode by operating directly in the pixel space and preserving the conditioning signal through topology-preserving mask injection applied throughout the decoding part, coupled with deterministic transport that does not repeatedly corrupt intermediate states with stochastic noise.
@@ 5 Conclusion
This paper introduced CrackSegFlow, a controllable flow-matching framework that renders realistic pavement images conditioned on an input crack mask while maintaining tight mask–image alignment. CrackSegFlow consists of two rendering modules—a topology-preserving mask injection module and a boundary-gated rendering module—that jointly enforce structural fidelity (connectivity and thin-structure continuity) and suppress mask-boundary artifacts, enabling the generated image to accurately match the provided mask under diverse pavement appearances. Because CrackSegFlow is explicitly mask-conditioned, the conditioning mask can be either real or synthetic, providing a unified mechanism for controllable data synthesis.
@@
To evaluate CrackSegFlow under a strong and consistent segmentation backbone, we developed a U-Net–style segmentor with a Transformer encoder (MiT-B4) together with a hybrid BCE + focal Tversky loss. While preliminary comparisons indicated that this backbone outperforms representative hybrid CNN–Transformer baselines, the main experiments did not aim to introduce a new segmentation architecture. Instead, the focus remained on how controllable flow-matching synthesis improves (i) in-domain performance, (ii) cross-domain transfer, and (iii) practical scalability relative to diffusion-based augmentation. The main conclusions of this study are as follows:
@@
In future work, we plan to extend the framework established in this paper to additional infrastructure distresses (e.g., potholes and surface spalling) and to study multi-distress controllable synthesis for improved field robustness. Moreover, while the main focus in this paper is controllable synthesis rather than segmentation-model development (and our U-MiT backbone provides a solid evaluation baseline), we note that in ongoing work we are developing a flow-matching-based crack segmentation model that shows strong promise on thin-structure benchmarks.

### v2

# CrackSegFlow: Controllable Flow Matching Synthesis for Generalizable Crack Segmentation with a 50K Image–Mask Benchmark

@@ Abstract
Automated crack segmentation is essential for condition assessment, yet deployment is limited by scarce pixel-level labels and domain shift. We present CrackSegFlow, a controllable Flow Matching synthesis framework that generates crack images conditioned on binary masks with mask–image alignment. The renderer combines topology-preserving mask injection with edge gating to maintain thin-structure continuity and suppress false positives. A class-conditional Flow Matching mask model synthesizes masks with control over crack coverage, enabling balanced, topology-diverse data without manual annotation. We inject masks into crack-free backgrounds to diversify illumination and reduce false positives. On five datasets with a CNN–Transformer backbone, incorporating synthesized pairs improves in-domain performance by 5.37 mIoU and 5.13 F1, and target-guided cross-domain synthesis yields gains of 13.12 mIoU and 14.82 F1 using target mask statistics. We also release CSF-50K, 50,000 image–mask pairs for benchmarking.
@@ 1 Introduction
    We introduce CrackSegFlow, a controllable Flow Matching (FM) synthesis framework that generates paired crack images and pixel-accurate masks with strict mask–image alignment. The image renderer integrates topology-preserving mask injection and boundary-gated modulation to preserve sub-pixel continuity of thin cracks and suppress texture-driven false positives under deterministic ODE sampling.
@@ 2 Related Works
Flow Matching (FM) Lipman et al. [2022] and Rectified Flow Liu et al. [2022] generate by learning a time-dependent vector field that transports a simple base distribution to data along a prescribed path, enabling deterministic ODE sampling. Training regresses closed-form pairwise velocities along the path, which yields stable objectives and few-step sampling. In dense prediction, two directions have emerged. First, unified Rectified Flow models bind image and mask spaces so a single network can both synthesize images and predict segmentation maps Wang et al. [2024a], Caetano and others [2025]. Second, segmentor-centric FM variants predict masks from images using flow-based objectives Bogensperger et al. [2025]. These works demonstrate FM’s viability but differ from an external, controllable augmentation engine; prior flow-based designs typically do not parameterize crack topology (e.g., explicit sparsity classes) nor guarantee geometry-consistent image rendering from masks at inference without additional labels.
@@ 3.2 Preliminaries
3.3 Semantic Flow Matching Model
@@ 3.5 Rectified Flow for Background-Guided Crack Injection
Given a new background $b$ and mask $M$, we solve the deterministic ODE $\frac{dx}{d\tau}=v_{\theta}(x,\tau,M),\quad x(0)=b,\ \tau\in[0,1],$ with a fixed-step Euler integrator (e.g., $K$ uniform steps). The terminal state $x(1)$ preserves the background context while injecting the requested crack topology specified by $M$. This complements our standard FM formulation by using the rectified schedule to “straighten” transport from background to target, yielding noise-free, deterministic sampling.
@@ 4.2.4 Synthesis protocol
As shown in  Fig.˜9, CrackSegFlow generates images that are precisely aligned with the conditioning masks. To maximize mask diversity, we first sample semantic masks using the mask generator and then render images conditioned on these masks. As depicted in  Fig.˜10, the resulting images are novel in crack geometry and texture relative to the originals. Table˜3 summarizes generative quality for both modalities. Across datasets, Flow Matching achieves low FID/KID, with masks consistently scoring best (reflecting sharp topology), and images remaining competitive. These scores corroborate the visual fidelity in Figs. 910 support the downstream gains, where higher-quality pairs correlate with improved mIoU/F1 after augmentation across all datasets.
@@
Table 3: Quality of Flow Matching synthesis measured by FID/KID (lower is better) for both mask and image modalities.
@@ 4.5 Comparison to diffusion-based semantic synthesis
Latent-diffusion pipelines, including LDM Rombach et al. [2022] and its conditioned variants such as ControlNet Zhang et al. [2023] and T2I-Adapter Mou et al. [2024], inject mask information through concatenation or auxiliary conditioning branches in the latent U-Net. While effective for general semantic layouts, this design can be sub-optimal for geometry-sensitive structures such as hairline cracks. In particular, repeated normalization and feature mixing can attenuate the conditioning signal at fine scales, leading to blurred filaments, weakened connectivity at junctions, and occasional mask–image drift Lei et al. [2025]. Fig.˜16 illustrates this limitation; latent-diffusion outputs often blur thin cracks or misalign them with the conditioning mask, whereas the FM formulation replaces stochastic noise injection with a continuous, condition-aware transport that preserves topology and alignment. CrackSegFlow avoids this failure mode by operating directly in the pixel space and preserving the conditioning signal through topology-preserving mask injection applied throughout the decoding part, coupled with deterministic transport that does not repeatedly corrupt intermediate states with stochastic noise.
@@ 5 Conclusion
This paper introduced CrackSegFlow, a controllable Flow Matching framework that renders realistic pavement images conditioned on an input crack mask while maintaining tight mask–image alignment. CrackSegFlow consists of two rendering modules—a topology-preserving mask injection module and a boundary-gated rendering module—that jointly enforce structural fidelity (connectivity and thin-structure continuity) and suppress mask-boundary artifacts, enabling the generated image to accurately match the provided mask under diverse pavement appearances. Because CrackSegFlow is explicitly mask-conditioned, the conditioning mask can be either real or synthetic, providing a unified mechanism for controllable data synthesis.
@@
To evaluate CrackSegFlow under a strong and consistent segmentation backbone, we developed a U-Net–style segmentor with a Transformer encoder (MiT-B4) together with a hybrid BCE + focal Tversky loss. While preliminary comparisons indicated that this backbone outperforms representative hybrid CNN–Transformer baselines, the main experiments did not aim to introduce a new segmentation architecture. Instead, the focus remained on how controllable Flow Matching synthesis improves (i) in-domain performance, (ii) cross-domain transfer, and (iii) practical scalability relative to diffusion-based augmentation. The main conclusions of this study are as follows:
@@
In future work, we plan to extend the framework established in this paper to additional infrastructure distresses (e.g., potholes and surface spalling) and to study multi-distress controllable synthesis for improved field robustness. Moreover, while the main focus in this paper is controllable synthesis rather than segmentation-model development (and our U-MiT backbone provides a solid evaluation baseline), we note that in ongoing work we are developing an FM-based crack segmentation model that shows strong promise on thin-structure benchmarks.
Acknowledgment
The first author gratefully acknowledges support from the University of Illinois Urbana-Champaign Graduate College through the Dissertation Completion Fellowship for the 2025–2026 academic year, awarded as the Department of Civil and Environmental Engineering recipient. The first author also thanks Professor Svetlana Lazebnik for helpful suggestions and feedback that improved the presentation of this work. This work used Delta GPU at the National Center for Supercomputing Applications (NCSA) through allocation CIS240456 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.